A cup of water is cooling. Its initial temperature is 100π‘œπΆ. After 3 minutes its temperature is 80π‘œπΆ. The temperature 𝑇 of the water, measured in ℃, is modelled by 𝑑𝑇 𝑑𝑑 = βˆ’π‘˜(𝑇 βˆ’ 25) where 𝑑 is the time elapsed in minutes. (i) Show that 𝑇 βˆ’ 𝐴𝑒 βˆ’π‘˜π‘‘ βˆ’ 25 = 0, where 𝐴 and π‘˜ are appropriate constants, (ii) Find the temperature of the water after 5 minutes.


1
Expert's answer
2021-10-12T10:10:59-0400

(i) Let "T(t)=" the temperature of a cup of waterΒ  at time "t."

The differential equation involving "T"


"\dfrac{dT}{dt}=-k(T-1000), k>0""\dfrac{dT}{T-25}=-kdt"

Integrate


"\int \dfrac{dT}{T-250}=-\int kdt"

"\ln(|T-25|)=-kt+\ln A"

"T-25=Ae^{-kt}"

"T-Ae^{-kt}-25=0"

(ii)

Given "T(0)=100\degree C, T(3)=80\degree C"

"100=25+Ae^{-k(0)}=>A=75"

"T(t)=25+75e^{-kt}"

"80=25+75e^{-k(3)}"

"e^{3k}=\dfrac{75}{55}"

"3k=\ln (\dfrac{15}{11})"

"k=\dfrac{1}{3}\ln (\dfrac{15}{11})"

"T(t)=25+75(\dfrac{11}{15})^{t/3}"


"T(5)=25+75(\dfrac{11}{15})^{5/3}"

"T(5)\approx69.7\degree C"